OpenStudy (anonymous):

P(t)=22cos(2.5piX)+98 The period of this function is the time between heartbeats. What is Dr. Evil's pulse rate (heartbeats per minute)?

OpenStudy (anonymous):

do you simply plug in 60 for t?

OpenStudy (anonymous):

p(t)=blood pressure

OpenStudy (jdoe0001):

\(\bf p(t)=22cos\left({\color{blue}{ \frac{5\pi}{2}}}x\right)\qquad period\implies \cfrac{2\pi}{{\color{blue}{ \frac{5\pi}{2}}}}\cdot 60mins \)

OpenStudy (jdoe0001):

woops, ahemm 60secs that is

OpenStudy (jdoe0001):

though there's an assumption that the rate given is per hearbeat per second

OpenStudy (jdoe0001):

well, I guess if it were per minute Mr. Evil would be dead =)

OpenStudy (anonymous):

AGREED.

OpenStudy (anonymous):

so it would be 48 with the equation that you have shown...is that possible? 48 beats per minute?

OpenStudy (jdoe0001):

I'd think so, yes notice the heart beats around once per sec or thereabouts

OpenStudy (anonymous):

i just dont understand the concept of the question. if the wquation is for blood pressure....how do blood pressure and pluse relate. i need an equation relating those.

OpenStudy (jdoe0001):

hmm dunno those ones....

OpenStudy (anonymous):

how about...if there are .8 seconds between heartbeats, then wouldnt we do 60/.8?

OpenStudy (anonymous):

that would be 75.